OpenStudy (anonymous):

∫_0^4▒(4e^5+ 3x)dx evaluate

OpenStudy (anonymous):

\[∫_0^4(4e^5+ 3x)dx \]

OpenStudy (anonymous):

again!!

OpenStudy (anonymous):

yes didnt understand it sorry!

OpenStudy (anonymous):

do you know how to find the anti derivative or are you supposed to use geometry?

OpenStudy (anonymous):

anti deri

OpenStudy (anonymous):

ok then this is straight forward application of power rule backwards the anti derivative of \[x^n\] is \[\frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

i think combination of both though anti + geometry

OpenStudy (anonymous):

is that possible?

OpenStudy (anonymous):

so the anti derivative of \[3x\] is \[\frac{3x^2}{2}\] and the anti derivative of \[4e^5\] is \[4e^5x\] is this ok?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so anti derivative of \[4e^5+3x\] is \[4e^5x+\frac{3x^2}{2}\]

OpenStudy (anonymous):

so we are almost done. last job is to replace x by 4

OpenStudy (anonymous):

\[∫_0^4(4e^5+ 3x)dx \]

OpenStudy (anonymous):

you get \[4e^5\times 4+\frac{3\times 4^2}{2}\] \[16e^5+24\]

OpenStudy (anonymous):

then replace x by 0. you get 0

OpenStudy (anonymous):

so the "final answer" is \[16e^5+24\] done

OpenStudy (anonymous):

plug in 0 to original form?

OpenStudy (anonymous):

oh no, plug 4 into anti- derivative, plug 0 into anti derivative and subtract

OpenStudy (anonymous):

wish you can write out everything man in one piece i am confused

OpenStudy (anonymous):

you are using this" \[\int_a^b f(x)dx =F(b)-F(a)\] where F is an anti derivative of f

OpenStudy (anonymous):

i can teach myself then

OpenStudy (anonymous):

k that makes sense

OpenStudy (anonymous):

ok here goes.

OpenStudy (anonymous):

\[f(x)=4e^5+3x\] \[F(x)=4e^5x +\frac{3x^2}{2}\] \[\int_0^4 (4e^5+3x) dx= F(4)-F(0)\] \[=4e^5\times 4+\frac{3\times 4^2}{2}-(4e^5\times 0+\frac{3\times 0^2}{2})\] \[=16e^5+24-0=16e^5+24\]

OpenStudy (anonymous):

all details there

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw